OpenStudy (anonymous):

how do yo simplify the ratio 14 to 22

OpenStudy (mertsj):

reduce the fraction

OpenStudy (dumbcow):

reduce by finding like factors 14/22 = 7*2/11*2 = 7/11

OpenStudy (mertsj):

\[\frac{14}{22}=\frac{7}{11}\]

OpenStudy (idily101):

7/11 because you divide the 14 and 22 by 2 to reduce